# Variational Bayes Gaussian Splatting

### v1

@@ Abstract
Recently, 3D Gaussian Splatting has emerged as a promising approach for modeling 3D scenes using mixtures of Gaussians. The predominant optimization method for these models relies on backpropagating gradients through a differentiable rendering pipeline, which struggles with catastrophic forgetting when dealing with continuous streams of data. To address this limitation, we propose Variational Bayes Gaussian Splatting (VBGS), a novel approach that frames training a Gaussian splat as variational inference over model parameters. By leveraging the conjugacy properties of multivariate Gaussians, we derive a closed-form variational update rule, allowing efficient updates from partial, sequential observations without the need for replay buffers. Our experiments show that VBGS not only matches state-of-the-art performance on static datasets, but also enables continual learning from sequentially streamed 2D and 3D data, drastically improving performance in this setting.
@@ 1 Introduction
Representing 3D scene information is a long-standing challenge for robotics and computer vision (Özyeşil et al., 2017). A recent breakthrough in this domain relies on representing the scene as a radiance field, i.e., using neural radiance fields (Mildenhall et al., 2020). Recently, Gaussian Splatting (Kerbl et al., 2023) has demonstrated the effectiveness of mixture models for 3D scene representation, as highlighted by a surge of subsequent research (see Chen and Wang (2024) for a survey). This approach leverages the ability of Gaussians to represent physical space as a collection of ellipsoids.
The dominant method for optimizing these models involves backpropagating gradients through a differentiable renderer with respect to the parameters of the mixture model. However, in many real-world applications, such as autonomous navigation, data is continuously streamed and must be processed sequentially. Traditional backpropagation-based methods are prone to catastrophic forgetting (French, 1999), leading to performance degradation as new data overwrites old knowledge. To mitigate this, replay buffers are often employed to retain and retrain on older data (Matsuki et al., 2024), which can be computationally expensive and memory intensive.
In this paper, we propose Variational Bayes Gaussian Splatting (VBGS), casting Gaussian Splatting as a variational inference problem over the parameters of a generative mixture model, enabling a closed-form update rule (Blei et al., 2017). Our approach inherently supports continual learning through iterative updates that are naturally accumulative, eliminating the need for replay buffers.
Our contributions are summarized as follows:
  • We propose a generative model of mixtures of Gaussians for representing spatial 2D (image) and 3D (point cloud) data and derive closed-form variational Bayes update rules.
  • We show that VBGS achieves a good model fit with a single parameter update step and enables continual learning without catastrophic forgetting.
  • We demonstrate and benchmark our approach on both 2D image (TinyImageNet (Le and Yang, 2015)) and 3D point cloud (Blender 3D models (Mildenhall et al., 2020) and Habitat scenes (Savva et al., 2019)) datasets.
@@ 2 Related Work
While our method also represents scenes as a mixture of Gaussians, it diverges from 3DGS by modeling a distribution over all features and parameters. Rather than relying on gradient-based optimization through differentiable rendering, we frame the learning process as variational inference. This approach allows us to perform continual learning through closed-form updates, enabling more efficient and robust handling of sequential data.
@@
Continual learning aims to develop models that can adapt to new tasks from a continuous datastream without forgetting previously learned knowledge (Wang et al., 2024). Backpropagation-based methods, such as 3DGS, face significant challenges in this area due to catastrophic forgetting (French, 1999), where the model’s performance on prior data deteriorates as it trains on new. This issue is particularly evident in real-world scenarios, such as localization and mapping, where data is streamed (Matsuki et al., 2024; Keetha et al., 2024). The common mitigation strategy involves maintaining a replay buffer of frames to revisit past data during updates (Sucar et al., 2021; Fu et al., 2024).
@@ 3 Method
where ${\theta}$ are the natural parameters of the distribution, $T({x})$ is the sufficient statistic, $A({\theta})$ is the log partition function, and $\phi({x})$ is the measure function. This likelihood has a conjugate prior of the following form:
@@ 3.1 The generative model
${s}$${c}$$z$$\Sigma_{{s},k}$$\mu_{{s},k}$$\mu_{{c},k}$$\Sigma_{{c},k}$$\pi$$N$$K$
Figure 1: The Generative Model: There are $K$ components in the mixture model, and $N$ observed data points, which are composed of a spatial vector ${s}$, and feature vector ${c}$. The parameters of the distributions of component $k$ ($\mu_{k}$, $\Sigma_{k}$) that generate ${s}$ and ${c}$ are random variables, which influence ${s}$ and ${c}$ respectively. $z$ is the associated mixture component for a given data point, dependent on the categorical parameters $\pi$. White and gray circles denote unobserved and observed variables respectively.
The generative model considered is a mixture model with $K$ components, each characterized by two conditionally independent modalities: spatial position (${s}$) and color (${c}$). For 2D images, ${s}$ represents the pixel location in terms of row and column coordinates (${s}\in\mathbb{R}^{2}$), while for 3D data, it corresponds to Cartesian coordinates (${s}\in\mathbb{R}^{3}$). The color component represents RGB values (${c}\in\mathbb{R}^{3}$) for both 2D and 3D data. Both ${s}_{k}$, and ${c}_{k}$ are modeled as multivariate Normal distributions with parameters $(\mu_{{s},k},\Sigma_{{s},k})$, and $(\mu_{{c},k},\Sigma_{{c},k})$ respectively, and the components have a mixture weight $z$. The full generative model is visualized as a Bayesian network in Figure 1 and is factorized as:
@@
To estimate the parameters, we infer their posterior distribution. However, as computing this posterior is intractable, we resort to variational inference (Jordan et al., 1998). We use a mean-field approximation to make inference tractable, which assumes that the variational posterior factorizes across the latent variables. This factorization allows for efficient coordinate ascent updates for each variable separately. Specifically, the variational distribution $q$ is decomposed as:
@@
We select the distributions of the approximate posteriors to be from the same family as their corresponding priors. For the color variable ${c}$, we model the mean by a Normal distribution but keep the covariance fixed using a Delta distribution. This assures that the mixture components commit to a particular color and do not blend multiple neighboring colors in a single component.
@@ 3.2 Coordinate ascent variational inference
For the approximate posterior $q(\mu_{{c},k},\Sigma_{{c},k})$ over the color likelihood, the sufficient statistics are again given by: $T({c}_{n})=({c}_{n},{c}_{n}\cdot{c}_{n}^{T})$. The prior is parameterized similarly as the NIW over ${s}$: $\eta_{0,{c}}=(\kappa_{0,{c}}\cdot m_{0,{c}},V_{0,{c}}+\kappa_{0,{c}}\cdot m_{0,{c}}\cdot m_{0,{c}}^{T})$ and $\nu_{0,{c}}=(\kappa_{0,{c}},n_{0,{c}}+D_{{c}}+1)$. However, as we model the prior over $\Sigma_{k,{c}}$ as a delta distribution, we keep the values for $n_{k,{c}}$ and $V_{k,{c}}$ fixed.
@@ 3.3 Continual updates
The proposed method supports continual learning because the parameters of each component are updated by aggregating the prior’s natural parameters with the data’s sufficient statistics via assignments $q(z)$. This iterative update process is order-invariant, allowing the model to adapt without forgetting past knowledge. Components without recent data assignments revert to their prior values, preserving flexibility in the model. Crucially, assignments $q(z)$ are always computed with respect to the initial posterior over parameters $q(\mu_{{s}},\Sigma_{{s}},\mu_{{c}},\Sigma_{{c}},\pi)$. This ensures that components without prior assignments can still be used to model the data.
@@
Hence, we can write an iterative update rule for the natural parameters at timestep $t$ as a function of the natural parameters at $t-1$:
@@ 4 Results
VBGS (Ours): We consider a variant of VBGS where the means of the initial posteriors ($q(\mu_{{s}},\Sigma_{{s}},\mu_{{c}},\Sigma_{{c}})$) are initialized on sampled points from the normalized data set (Data Init), as well as randomly initialized ($m_{k,{s}}\sim\mathcal{U}[-1,1]$, $m_{k,{c}}\sim\delta(0)$) (Random Init). For models with random initialization, data normalization is performed using estimated statistics; see Appendix C for further details.
Gradient: In 3DGS (Kerbl et al., 2023), the parameters are directly optimized using stochastic gradient descent on a weighted image reconstruction loss ($(1-\lambda)\cdot\text{MSE}+\lambda\cdot\text{SSIM}$). In the case of image data, $\lambda$ is set to 00; in the case of 3D data $\lambda$, it is set to $0.2$. We use spherical harmonics with no degrees of freedom, i.e., the specular reflections are not modeled. In order to be able to compare performance w.r.t. model size, we also do not do densification or shrinking (Kerbl et al., 2023) (for these results see Appendix D.2). When optimizing for images, we use a fixed camera pose at identity and keep the z-coordinate of the Gaussians fixed at a value of 1. Similar to the VBGS approach, we also consider a variant where the means of the Gaussian components are randomly initialized (Random Init), and on sampled points from the dataset (Data Init).
@@ 4.1 Images
We first evaluate performance on the Tiny ImageNet test set consisting of 10k images. Reconstruction accuracy, measured in PSNR (dB), is evaluated across varying numbers of components (Figure 2). Our results indicate that VBGS achieves reconstruction errors comparable to the gradient-based approach. We then evaluate the models in a continual learning setting, where data patches are sequentially observed and processed (Figure 2). While VBGS maintains consistent reconstruction quality across all observed patches, the gradient-based method disproportionately focuses on the most recent patch. The PSNR over timesteps is shown in Figure 3. VBGS converges to the same accuracy as training on the static dataset, whereas Gradient degrades due to catastrophic forgetting.
We further compared the computational efficiency by measuring the wall-clock time required for the Gradient approach to reach the performance level of VBGS after a single update step. This was computed over all images of the Tiny ImageNet validation set. We observe that VBGS is significantly (t-test, $p=0$) faster in wall clock time ($0.03\pm 0.03$ seconds) compared to Gradient ($0.05\pm 0.02$ seconds).
To assess performance in a continual learning setting, we divide each image into 8x8 patches, feeding the model one patch at a time (Figure 2). VBGS performs a single update per patch, whereas the Gradient approach performs 100 training steps per patch with a learning rate of 0.1. Figure 3 shows how the reconstruction PSNR evolves at each time step over the tiny ImageNet test set for the model with capacity 10K components. It’s important to note that the reconstruction error of VBGS after observing all the patches converges to the same value as observing the whole image at once, as the inferred posterior ends up being identical. Even though the gradient approach, when observing all data at once, achieves much higher PSNR, when the data is fed in continuously, it does not achieve that performance in the continual setting, as it always focuses on the latest observed patch (see Figure 2).
@@ 4.2 Blender 3D Objects
Next, we evaluate VBGS on 3D objects from the Blender dataset (Mildenhall et al., 2020). Model parameters are inferred using 200 frames which include depth information. VBGS is trained on the 3D point cloud, which is acquired by transforming the RGBD frame to a shared reference frame. In contrast, the gradient-based approach is optimized using multi-view image reconstruction. We evaluate reconstruction performance on 100 frames from the validation set. The results, measured as PSNR, for a model with a capacity of 100K components are shown in Table 1. Our results show improved performance for both methods when components are initialized from data rather than randomly. VBGS achieves performance similar to the gradient-based approach under Data Init conditions. In random initialization, our approach outperforms the gradient-based approach except for the “ship” object. For a more in-depth analysis where we vary the number of components, see Appendix D.1.
@@
VBGS (Data ${}^{\text{(Data Init)}}$ $22.82$0.94 $\mathbf{19.50}$0.47 $\mathbf{22.06}$0.79 $\mathbf{23.62}$1.23 $\mathbf{22.53}$0.89 $\mathbf{20.55}$1.64 $24.78$0.60 $21.23$0.64
VBGS (Random ${}^{\text{(Random Init)}}$ $21.35$0.69 $18.71$0.43 $21.49$0.75 $22.24$0.97 $20.59$0.86 $20.47$1.36 $24.42$0.58 $20.80$0.90
Gradient (Data ${}^{\text{(Data Init)}}$ $\mathbf{22.98}$1.23 $19.05$0.57 $21.08$0.92 $21.47$1.67 $19.97$2.24 $20.53$1.64 $\mathbf{25.25}$0.90 $\mathbf{23.55}$1.20
Gradient (Random ${}^{\text{(Random Init)}}$ $20.59$0.85 $15.04$1.13 $19.41$0.90 $19.81$1.86 $19.10$1.02 $16.11$1.45 $23.03$0.72 $21.15$0.82
@@
4.3 Habitat Rooms
Finally, we evaluate on a dataset of 3D rooms, which is relevant for e.g. autonomous navigation for robotics in unseen environments. We consider three rooms from the Habitat test suite (“Van Gogh Room”, “Apartment 1” and “Skokloster Castle”) (Savva et al., 2019). The model parameters are inferred using 200 randomly sampled views from each room using the preprocessing pipeline of (Wang et al., 2023). On top of the models from the previous sections, we now also evaluate the model with component reassignments as detailed in Section 3.4.

### v2

@@ Abstract
Recently, 3D Gaussian Splatting has emerged as a promising approach for modeling 3D scenes using mixtures of Gaussians. The predominant optimization method for these models relies on backpropagating gradients through a differentiable rendering pipeline, which struggles with catastrophic forgetting when dealing with continuous streams of data. To address this limitation, we propose Variational Bayes Gaussian Splatting (VBGS), a novel approach that frames training a Gaussian splat as variational inference over model parameters. By leveraging the conjugacy properties of multivariate Gaussians, we derive a closed-form variational update rule, allowing efficient updates from partial, sequential observations without the need for replay buffers. Our experiments show that VBGS enables continual learning from sequentially streamed 2D and 3D data, drastically improving performance over the baseline.
@@ 1 Introduction
Representing 3D scene information is a long-standing challenge for robotics and computer vision (Özyeşil et al., 2017). A recent breakthrough in this domain relies on representing the scene as a radiance field, i.e., using Neural Radiance Fields (NeRF) (Mildenhall et al., 2020). Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has demonstrated the effectiveness of mixture models for 3D scene representation, as highlighted by a surge of subsequent research (see Chen and Wang (2024) for a survey). This approach leverages the ability of Gaussians to represent physical space as a collection of ellipsoids. The dominant method for optimizing these models involves backpropagating gradients through a differentiable renderer with respect to the parameters of the mixture model.
While effective, 3DGS – and gradient-based methods more broadly – faces critical challenges in continual learning scenarios. Many real-world applications, such as autonomous navigation, simultaneous localization and mapping (SLAM), and robotics involve data that is continuously streamed and must be processed sequentially. In such settings, gradient-based optimization methods are prone to catastrophic forgetting (French, 1999), leading to performance degradation as new data overwrites old knowledge. To mitigate this, replay buffers are often employed to retain and retrain on older data (Matsuki et al., 2024), which can be computationally expensive and memory intensive.
To overcome these challenges, we propose Variational Bayes Gaussian Splatting (VBGS), casting Gaussian Splatting as a variational inference problem over the parameters of a generative mixture model, enabling a closed-form update rule (Blei et al., 2017). Our approach inherently supports continual learning through iterative updates that are naturally accumulative, eliminating the need for replay buffers.
Our contributions are summarized as follows: (i) We propose a generative model of mixtures of Gaussians for representing spatial 2D (image) and 3D (point cloud) data and derive closed-form variational Bayes update rules. (ii) We show that VBGS enables continual learning without catastrophic forgetting, and VBGS models achieve a good model fit with only a single parameter update step. (iii) We demonstrate and benchmark our approach on both 2D image (TinyImageNet (Le and Yang, 2015)) and 3D point cloud (Blender 3D models (Mildenhall et al., 2020) and Habitat scenes (Savva et al., 2019)) datasets.
@@ 2 Related Work
While our method also represents scenes as a mixture of Gaussians, it diverges from 3DGS by modeling a distribution over all features and parameters similar to (Kheradmand et al., 2025). However, in contrast to (Kheradmand et al., 2025), our methods does not rely on gradient-based optimization through differentiable rendering, we frame the learning process as variational inference. This approach allows us to perform continual learning through closed-form updates, enabling more efficient and robust handling of sequential data.
@@
Continual learning aims to develop models that can adapt to new tasks from a continuous datastream without forgetting previously learned knowledge (Wang et al., 2024). Backpropagation-based methods, such as 3DGS, face significant challenges in this area due to catastrophic forgetting (French, 1999), where the model’s performance on prior data deteriorates as it trains on new. This issue is particularly evident in real-world scenarios, such as localization and mapping, where data is streamed (Matsuki et al., 2024; Keetha et al., 2024). The common mitigation strategy involves maintaining a replay buffer of frames to revisit past data during updates (Sucar et al., 2021; Fu et al., 2024; Zhipeng Cai, 2023).
@@ 3 Method
where $\theta$} are the natural parameters of the distribution, T(x)T({x}) is the sufficient statistic, $A({\theta})$ is the log partition function, and $\phi({x})$ is the measure function. This likelihood has a conjugate prior of the following form:
@@ 3.1 The generative model
[…]
Figure 1: The Generative Model: Mixture model with $K$ components, and NN observed data points, which are composed of a spatial vector s{s}, and feature vector c{c}. The parameters of the distributions of component $k$ ($\mu_{k}$, $\Sigma_{k}$) that generate s{s} and c{c} are random variables, which influence s{s} and c{c} respectively. $z$ is the associated mixture component for a given data point, dependent on the categorical parameters $\pi$. White and gray circles denote unobserved and observed variables.
The generative model considered is a mixture model with $K$ components, each characterized by two conditionally independent modalities: spatial position (s{s}) and color (c{c}). For 2D images, s{s} represents the pixel location in terms of row and column coordinates (${s}\in\mathbb{R}^{2}$), while for 3D data, it corresponds to Cartesian coordinates (${s}\in\mathbb{R}^{3}$). The color component represents RGB values (${c}\in\mathbb{R}^{3}$) for both 2D and 3D data. Both ${s}_{n}$, and ${c}_{n}$ are modeled as multivariate Normal distributions with parameters $(\mu_{{s},k},\Sigma_{{s},k})$, and $(\mu_{{c},k},\Sigma_{{c},k})$ respectively, and the components have a mixture weight $z$. The full generative model is visualized as a Bayesian network in Figure 1 and is factorized as:
@@
To estimate the parameters, we infer their posterior distribution. However, as computing this posterior is intractable, we resort to variational inference (Jordan et al., 1998). We use a mean-field approximation to make inference tractable, which assumes that the variational posterior factorizes across the latent variables, as otherwise the interaction between the various Gaussian components makes the inference process grow combinatorially. This factorization allows for efficient coordinate ascent updates for each variable separately. Specifically, the variational distribution $q$ is decomposed as:
@@
We select the distributions of the approximate posteriors to be from the same family as their corresponding priors. For the color variable c{c}, we model the mean by a Normal distribution but keep the covariance fixed using a Delta distribution. This assures that the mixture components commit to a particular color and do not blend multiple neighboring colors in a single component.
@@ 3.2 Coordinate ascent variational inference
For the approximate posterior $q(\mu_{{c},k},\Sigma_{{c},k})$ over the color likelihood, the sufficient statistics are again given by: $T({c}_{n})=({c}_{n},{c}_{n}\cdot{c}_{n}^{T})$. The prior is parameterized similarly as the NIW over s{s}: $\eta_{0,{c}}=(\kappa_{0,{c}}\cdot m_{0,{c}},V_{0,{c}}+\kappa_{0,{c}}\cdot m_{0,{c}}\cdot m_{0,{c}}^{T})$ and $\nu_{0,{c}}=(\kappa_{0,{c}},n_{0,{c}}+D_{{c}}+1)$. However, as we model the prior over $\Sigma_{k,{c}}$ as a delta distribution, we keep the values for $n_{k,{c}}$ and $V_{k,{c}}$ fixed.
@@ 3.3 Continual updates
A major benefit of VBGS is that it enables continuous learning because the parameters of each component are updated by aggregating the prior’s natural parameters with the data’s sufficient statistics through its assignments $q(z)$. This iterative update process is order-invariant, allowing the model to adapt without forgetting past knowledge. Components without recent data assignments revert to their prior values, preserving flexibility in the model. Crucially, assignments $q(z)$ are always computed with respect to the initial posterior over parameters $q(\mu_{{s}},\Sigma_{{s}},\mu_{{c}},\Sigma_{{c}},\pi)$. This ensures that components without prior assignments can still be used to model the data.
@@
Hence, we can write an iterative update rule for the natural parameters at timestep $t$ as a function of the natural parameters at t1t-1:
@@ 4 Results
VBGS (Ours): We consider a variant of VBGS where the means of the initial posteriors ($q(\mu_{{s}},\Sigma_{{s}},\mu_{{c}},\Sigma_{{c}})$) are initialized on sampled points from the normalized data set (Data Init), as well as randomly initialized ($m_{k,{s}}\sim\mathcal{U}[-1,1]$, $m_{k,{c}}\sim\delta(0)$) (Random Init). For models with random initialization, data normalization is performed using estimated statistics; see Appendix C.
Gradient: In 3DGS (Kerbl et al., 2023), the parameters are directly optimized using stochastic gradient descent on a weighted image reconstruction loss ($(1-\lambda)\cdot\text{MSE}+\lambda\cdot\text{SSIM}$). In the case of image data, $\lambda$ is set to 0; in the case of 3D data $\lambda$, it is set to $0.2$. We use spherical harmonics with no degrees of freedom, i.e., the specular reflections are not modeled. In order to be able to compare performance w.r.t. model size, we also do not do densification or shrinking (Kerbl et al., 2023) (for these results see Appendix D.2). When optimizing for images, we use a fixed camera pose at identity and keep the z-coordinate of the Gaussians fixed at a value of 1. Similar to the VBGS approach, we also consider a variant where the means of the Gaussian components are randomly initialized (Random Init), and on sampled points from the dataset (Data Init).
@@ 4.1 Images
We first evaluate performance on the Tiny ImageNet test set consisting of 10k images. We measure reconstruction accuracy using the Peak Signal-to-Noise Ratio (PSNR) metric, which expresses pixel error on a logarithmic scale (dB) . Reconstruction accuracy is evaluated across varying numbers of components (Figure 2). Our results indicate that VBGS achieves reconstruction errors comparable to the gradient-based approach. We then evaluate the models in a continual learning setting, where data patches are sequentially observed and processed (Figure 2). While VBGS maintains consistent reconstruction quality across all observed patches, the gradient-based method disproportionately focuses on the most recent patch. The PSNR over timesteps is shown in Figure 3. VBGS converges to the same accuracy as training on the static dataset, whereas Gradient degrades due to catastrophic forgetting. We also compared the computational efficiency by measuring the wall-clock time required for the Gradient approach to reach the performance level of VBGS after a single update step using an NVIDIA GeForce RTX 4090. This was computed over all images of the Tiny ImageNet validation set. We observe that VBGS is significantly (t-test, $p=0$) faster in wall clock time ($0.03\pm 0.03$ s) compared to Gradient ($0.05\pm 0.02$ s).
To assess performance in a continual learning setting, we divide each image into 8x8 patches, feeding the model one patch at a time (Figure 2). VBGS performs a single update per patch, whereas the Gradient approach performs 100 training steps per patch with a learning rate of 0.1. Figure 3 shows how the reconstruction PSNR evolves at each time step over the tiny ImageNet test set for the model with capacity 10K components. It’s important to note that the reconstruction error of VBGS after observing all the patches converges to the same value as observing the whole image at once, as the inferred posterior ends up being identical. Even though the gradient approach achieves much higher PSNR when observing all data at once, it fails to achieve that performance in the continual setting, as it always focuses on the latest observed patch (see Figure 2).
@@ 4.2 Blender 3D Objects
Next, we evaluate VBGS on 3D objects from the Blender dataset (Mildenhall et al., 2020). Model parameters are inferred using 200 frames which include depth information. VBGS is trained on the 3D point cloud, which is acquired by transforming the RGBD frame to a shared reference frame. In contrast, the gradient-based approach is optimized using multi-view image reconstruction. We evaluate reconstruction performance on 100 frames from the validation set. The results, measured as PSNR, for a model with a capacity of 100K components are shown in Table 1. Our results show improved performance for both methods when components are initialized from data rather than randomly. VBGS achieves performance similar to the gradient-based approach under Data Init conditions. In random initialization, our approach outperforms the gradient-based approach except for the “ship” object. For a more in-depth analysis on the number of components, see Appendix D.1.
@@
VBGS $^{\text{(Data Init)}}$ $22.82$ ±0.94 $\mathbf{19.50}$ ±0.47 $\mathbf{22.06}$ ±0.79 $\mathbf{23.62}$ ±1.23 $\mathbf{22.53}$ ±0.89 $\mathbf{20.55}$ ±1.64 $24.78$ ±0.60 $21.23$ ±0.64
VBGS $^{\text{(Random Init)}}$ $21.35$ ±0.69 $18.71$ ±0.43 $21.49$ ±0.75 $22.24$ ±0.97 $20.59$ ±0.86 $20.47$ ±1.36 $24.42$ ±0.58 $20.80$ ±0.90
Gradient $^{\text{(Data Init)}}$ $\mathbf{22.98}$ ±1.23 $19.05$ ±0.57 $21.08$ ±0.92 $21.47$ ±1.67 $19.97$ ±2.24 $20.53$ ±1.64 $\mathbf{25.25}$ ±0.90 $\mathbf{23.55}$ ±1.20
Gradient $^{\text{(Random Init)}}$ $20.59$ ±0.85 $15.04$ ±1.13 $19.41$ ±0.90 $19.81$ ±1.86 $19.10$ ±1.02 $16.11$ ±1.45 $23.03$ ±0.72 $21.15$ ±0.82
@@
4.3 3D Rooms
Figure 5: Continual learning results for the Replica rooms dataset. (a) Evolution of the reconstruction performance, measured in PSNR (dB), after feeding in consecutive images of the room. The shaded area indicates the 95% confidence interval computed over the 100 frames from the validation set. (b) Qualitative result of reconstructions throughout the observed sequence on Room 0 of the Replica dataset. The time $t$ indicates the number of observed frames, while the frame index refers to the camera pose from which the model is rendered.
Finally, we evaluate the performance on the Replica dataset (Straub et al., 2019), which is often used to benchmark robotic SLAM systems – and hence a good candidate for continual learning. We train a model using a reassign fraction of $0.01$. Sequences in the Replica datasets consist of 2000 frames. We train our model on every 10th frame, starting by frame 0 (resulting in 200 train frames), and evaluate on every 20th frame, starting by frame 5 (resulting in 100 evaluation frames).
Figure 5 depicts the evolution of PSNR after observing multiple frames in a moving trajectory. The PSNR increases steadily as larger aspects of the room are observed (e.g. around step 75 in Figure 5, the camera turns around and observes the other side of the room). In the continual learning setting, the gradient baseline does not improve reconstruction quality. We also report the performance of the gradient baseline in an offline setting with a replay buffer. The discrepancy clearly shows the effect of catastrophic forgetting. Note that at first some parts of the room are not yet observed and it can not generate each view (e.g. frame 50 at t=0). However, for each of the datasets, after frame 100 the entire room has been observed and can be accurately reconstructed without sacrificing reconstruction performance on the first observed frame. For additional results, see Appendix D.4.